OpenStudy (anonymous):

Two vectors are given by vector a = 6i+8.7j and vector b = 3.2 i +8j, the component of a vector a along the direction of vector b?

OpenStudy (anonymous):

To find the component of any vector a along any vector b, you need to dot a with the unit vector along b.

OpenStudy (anonymous):

oh thank you :)

OpenStudy (anonymous):

my answer is 89, it's incorrect :( @mashy

OpenStudy (shamim):

The component of a along b is a*cos theta=?

OpenStudy (shamim):

We know a.b=a*b*cos theta a*cos theta=a.b/b=89/b=?

OpenStudy (shamim):

Feel free to ask for more clarification